AT WHAT ANGLE WITH THE HORIZONTAL SHOULD A BALL BE THRWN SO THAT THE RANGE R IS RELATED TO TIME OF FLIGHT ‘T’ AS R=5T^2

Grade:9

1 Answers

Arun
6 years ago
R=5T^2
(u^2*sin2 theta)/g=5*[(2*u*sin theta)/g]^2
(u^2*sin2 theta)/g=5*4*u^2*sin^ 2theta/g^2
sin2 theta*g=5*4*sin^2theta
10sin2 theta=20sin^2 theta (g=10)
sin2 theta=2sin^2 theta
2sin theta*cos theta=2sin^2 theta  (sin2 theta=2sin theta*cos theta)
sin theta*cos theta=sin^2 theta
cos theta = sin theta
cot theta = 1
i.e. theta= 45 degree